OpenStudy (anonymous):

5=(x/3) (1/8 - 1/7) is type of question how does it end up as 5=#x shoudn't it be 5=x/#

OpenStudy (anonymous):

5=(x/3)(7-8)/(7*8) x/3=-1*56 x=-168

OpenStudy (anonymous):

Did you subtract correctly? \[\frac{1}{8}- \frac{1}{7}=\frac{7-8}{7 \times 8}\]